. Pro-Sports, a sports equipment manufacturer, has a machine currently in use that was originally purchased two years ago for $80,000. The firm is depreciating the machine on a straight-line basis using a five-year recovery period. The present machine will last for the next five years or, once removal and clean-up costs are taken into consideration, it could be sold now for $50,000. Pro-Sports can buy a new machine today for a net price of $120,000 (including all installation costs). The proposed machine will be depreciated using a five-year straight-line depreciation period. If the firm acquires the new machine, there will be no change in its investment in net working capital. Profit before depreciation and taxes (PBDT) is expected to be $90,000 for each of next five years for the present machine, and $110,000 for each of next five years for the proposed machine. The corporate tax rate for the firm is 28%. Pro-Sports expects to be able to sell the proposed machine at the end of its five-year usable life for $20,000 (after paying removal and clean-up costs). Alternatively, the present machine is expected to net $5,000 on its sale at the end of the same period. Required: A. Calculate the initial investment associated with the replacement decision. [2 marks] B. Calculate the incremental operating cash flows associated with the proposed replacement decision. [4 marks] C. Calculate the terminal cash flow associated with the proposed replacement decision. [2 marks] D. Assuming Pro-Sports has a WACC of 15%, please make a recommendation to the management as to whether or not the old machine should be replaced. Show all calculations to support your recommendation.
Answer #1
A Initial investment with replacement decision
a Investment in Replacement machine $120,000
b Annualdepreciation of old machine $16,000 (80000/5)
c=b*2 Accumulated depreciation of old machine $32,000
d Book valueof the old machine $48,000 (80000-32000)
e Current Salvage value before tax $50,000
f=e-d Gain on salvage $2,000
g=f*28% Tax on gain $560
h=e-g After tax salvage value of existing machine $49,440
i=a-h Initial invetment with replacement decision $70,560
B Incremental Operating Cash flow
a Annual depreciation of existing machine $16,000
b Annual depreciation of Replacement machine $24,000 (120000/5)
c=b-a Incremental Depreciation $8,000
d=c*28% Depreiation Tax Shield(Incremental) $2,240
e PBDT fromexisting machine $90,000
f PBDT from replacement machine $110,000
g=f-e Incremental Profit before taxes $20,000
h=g*(1-0.28) Incremental Profit after taxes $14,400
i=d+h Incremental Operating Cash flow $16,640
C Terminal Cash Flow
a Salvage value from existing machine $5,000
b Salvage value from replacement machine $20,000
c=b-a Before tax terminal cash flow $15,000
d=c*(1-0.28) After tax terminalcash flow $10,800
D WACC=15%=0.15
Present Value of Cash Flow=(Cash flow)/((1+i)^N)
i=discount rate=WACC=0.15
N=year of cash flow
N YEAR 0 1 2 3 4 5
a Initial Cash Flow ($70,560)
b Annual Operating Cash Flow $16,640 $16,640 $16,640 $16,640 $16,640
c TerminalCash flow $10,800
d=a+b+c Net cash Flow ($70,560) $16,640 $16,640 $16,640 $16,640 $27,440 SUM
PV=d/(1.15^N) Present Value of net cash flow ($70,560) $14,470 $12,582 $10,941 $9,514 $13,643 ($9,411)
NPV=Sumof PV Net Present Value ($9,411)
Old machine should not be replace
Because PV is negative
The Replacement does not give required return
InternalRate of Return (IRR) 9.6% (Using IRR function of excelover Net Cash flow)